OpenStudy (anonymous):

Find the solution of the given initial value problem: y''+3y=0 y(0)= -2 y'(0)= 3

OpenStudy (anonymous):

Well I figured it out if anyone was interested: y=e^(mx) y'=me^(mx) y''=(m^2)e^(mx) Knowing this, just plug it back into the equation: (m^2)e^(mx)+3(e^(mx)) = 0 Divide by e^(mx) which gets you: m^2 + 3 = 0 Solve for m: m= sqrt(3) and m= -sqrt(3) General solution: y=C1e^sqrt(3) + C2e^-sqrt(3) Use initial values: -2 = C1 + C2 3 = sqrt(3)C1 - sqrt(3)C2 Solve for C1 and C2 by substitution and plug them back into the general solution!

OpenStudy (anonymous):

You made a slight mistake in solving \(m^2+3=0\). The equation has two roots\(m_1=\sqrt{3}i\) and \(m_2=-\sqrt{3}i\). The solution will be then \(y(x)=c_1\cos(\sqrt{3}x)+c_2\sin(\sqrt{3}x)\). You can now find the constants by pluggin the given initial values.

OpenStudy (anonymous):

If you're wondering how the solution came to be in terms of cosine and sine functions, then you should know that \(e^{ix}=\cos{x}+i\sin{x}\).